# Morse inequalities for noncompact manifolds

### v1

@@ 1. Introduction
Actually, boundedness of the gradient vector field $\nabla f$ is essential, and that of $f$ and its higher derivatives may be a technical condition. We note that the boundedness of a function is independent of the choice of a $G$-invariant metric on $M$ because $G$-invariant metrics on $M$ are mutually equivalent. By (1.2) below, we describe a rough configuration of critical points of a Morse function on $M$ by a function on $G$, which may be thought of as counting of critical points. In order to make this function controllable, we assume the following uniformness of Morse functions. We fix a $G$-invariant triangulation of $M$. Then as in Section 2, we can construct a fundamental domain $K$ as a finite subcomplex of $M$ such that
(1.1) $M=\bigcup_{g\in G}gK$
and $gK\cap hK$ has dimension at most $n-1$ for $g\neq h\in G$.
@@
A Morse function $f\colon M\to\mathbb{R}$ is uniform if there is $\epsilon>0$ satisfying that
  1. (1)
    any two critical points of $f$ are at least $2\epsilon$ distant, and
  2. (2)
    for each critical point $p$ of $f$, there is $g\in G$ such that the $\epsilon$-neighborhood of $p$ is contained in $gK$.
@@ A Morse function $f\colon M\to\mathbb{R}$ is uniform if there is $\epsilon>0$ satisfying that
Then this is a function describing a rough configuration of critical points of $f$ with index $k$, mentioned above. This way of counting of discrete points on $M$ was discovered in [12]. Observe that $c_{k}$ is a bounded function for each $k\geq 0$ whenever $f$ is uniform, because a fundamental domain $K$ is compact.
@@
We prove Theorem 1.3 by employing the Witten deformation by a Morse function. Then we need to disassemble the trace of an operator on the Hilbert space of $L^{2}$-sections of a vector bundle over $M$ with respect to the cover (1.1). In Section 2, we introduce a piecewise trace-class operator and its piecewise trace which is the desired disassembly of the usual (or Roe’s) trace. In Section 3, we introduce an operator whose kernel function has propagation subordinate to a Gaussian function centered at the diagonal set, which we call an operator with Gaussian propagation. We prove the trace property of the piecewise trace of a product of operators with Gaussian propagation. We also consider the functional calculus of a generalized Dirac operators such that the resulting operator has Gaussian propagation. In Section 4, we introduce a disassembly of the normalized Betti number of Roe [19, 20], which is related to the $L^{2}$-Betti number of $M$, with respect to the cover (1.1), and prove its invariance under a certain deformation, including the Witten deformation by a bounded function. In Section 5, we collect all results obtained so far, and apply the Witten deformation to prove Theorem 1.3 and its corollaries. The boundedness of a Morse function itself is used only in the proof of the invariance of the normalized Betti number, though the boundedness of the derivative is essential as mentioned above. Then we pose some question on the boundedness of a Morse function itself.
@@ 2.1. Fundamental domain
First of all, we recall from [12] the construction of a fundamental domain. Choose a $G$-invariant triangulation $L$ of $M$, that is, the lift of a triangulation of $M/G$. Then $L/G$ is a triangulation of $M/G$. For each $n$-simplex of $L/G$, we choose one of its lifts to $L$, and define the subcomplex $K$ of $L$, called a fundamental domain, by the union of those $n$-simplices of $L$. Remark that a fundamental domain $K$ needs not be connected. By definition, we have
$M=\bigcup_{g\in G}gK$
and $\dim(gK\cap hK)\leq n-1$ for $g\neq h\in G$, implying $\mathrm{Int}(gK)\cap\mathrm{Int}(hK)=\emptyset$ for $g\neq h$. In the sequel, we fix a fundamental domain $K$.
@@ 5.1. Strongly uniform Morse function
The key technical ingredient of Morse theory is the Morse lemma which states that near a critical point, a Morse function is identified with a nondegenerate quadratic function by choosing an appropriate coordinate. Such a coordinate is not generally a normal coordinate, which is inconvenient for our purpose. Then, instead of Morse lemma, we consider replacement of a given Morse function by a nice Morse function with the same critical points. We show that this is possible for a strongly uniform Morse function, and so we only need to prove Theorem 1.3 for such a nice Morse function.
@@
A bounded Morse function $f\colon M\to\mathbb{R}$ is called strongly uniform if it is uniform with respect to $\epsilon>0$ and satisfies
$\inf_{\lambda}|\lambda|=\delta_{1}\quad\text{and}\quad\inf_{x\in M-N(\operatorname{Crit}(f);\epsilon_{0})}|\nabla f(x)|\geq\delta_{2}$
for some $\delta_{1},\delta_{2}>0$, where $\lambda$ ranges over all eigenvalues of the Hessian $\nabla^{2}f(p)$ for all $p\in\operatorname{Crit}(f)$ and
@@
Let $g$ denote the fixed $G$-invariant metric on $M$. We introduce a new $G$-invariant metric on $M$. Let $L$ denote the fixed $G$-invariant triangulation of $M$, which is used to construct a fundamental domain $K$. Take any small $\epsilon>0$. We define a $G$-invariant metric $g(\epsilon)$ on $M$ such that $g(\epsilon)=g$ on $N(L_{n-1};\frac{\epsilon}{2})$ and $g(\epsilon)$ is flat on the complement of $N(L_{n-1};\epsilon)$, where $L_{n-1}$ denotes the $(n-1)$-skeleton of $L$. Indeed, such a $G$-invariant metric $g(\epsilon)$ exists. Let $f\colon M\to\mathbb{R}$ be a uniform Morse function with respect to $\epsilon>0$. Then for any critical point $p$ of $f$, $N(p;\epsilon)$ is in the complement of $N(L_{n-1};\epsilon)$, and so the metric $g(\epsilon)$ is flat on $N(p;\epsilon)$. We remark that all results in the previous sections hold for any $G$-invariant metric on $M$, hence for $g(\epsilon)$. As mentioned above, boundedness of a function is independent of the choice of a $G$-invariant metric on $M$. We will use the fixed metric $g$ for considering upper and lower bounds for a Morse function.
@@
Take any $p\in\operatorname{Crit}_{k}(\tilde{f})$. Then $p\in g(\mathrm{Int}(K))$ for some $g\in G$, so $g$ is unique. Let $\rho_{p}(x)=\tilde{\rho}(d(x,p))$. Recall that we are working with the metric $g(\epsilon)$ on $M$, which is flat on $N(p;\epsilon)$ for each $p\in\operatorname{Crit}(\tilde{f})$. Let $\lambda_{1},\ldots,\lambda_{n}$ be the eigenvalue of the Hessian $\nabla^{2}\tilde{f}(p)$, and let
@@
As in the proof of Proposition 2.12, we have $\mu(\mathcal{I})=0$. Then the statement follows from Theorem 1.3 and the positivity of $\mu$. ∎
@@
Then we get $\mu(c_{k})\geq b_{k}^{(2)}$ for any $G$-invariant mean $\mu\colon\ell^{\infty}(G)\to\mathbb{R}$. Hence by Proposition 5.8, $c_{k}(g)\neq 0$ for infinitely many $g\in G$. ∎
@@
5.3. Questions on boundedness

### v2

@@ 1. Introduction
Actually, boundedness of the gradient vector field $\nabla f$ is essential, and that of $f$ and its higher derivatives may be a technical condition. We note that the boundedness of a function is independent of the choice of a $G$-invariant metric on $M$ because $G$-invariant metrics on $M$ are mutually equivalent. By (1.2) below, we describe a rough configuration of critical points of a Morse function on $M$ by a function on $G$, which may be thought of as counting of critical points. In order to make this function controllable, we assume the following uniformness of Morse functions. We fix a $G$-invariant triangulation of $M$. Then as in Section 2, we can construct a fundamental domain $K$ such that the closure of $K$ is a finite subcomplex of $M$ and it satisfies
(1.1) $M=\coprod_{g\in G}gK.$
Here, we note that the closure of $K$ is a fundamental domain in the sense of [12].
@@
A Morse function $f\colon M\to\mathbb{R}$ is uniform if there is $\epsilon>0$ such that any two critical points of $f$ are at least $2\epsilon$ distant.
@@
Then this is a function describing a rough configuration of critical points of $f$ with index $k$, mentioned above. This way of counting of discrete points on $M$ was discovered in [12]. Observe that $c_{k}$ is a bounded function for each $k\geq 0$ whenever $f$ is uniform, because the closure of a fundamental domain $K$ is compact.
@@
We prove Theorem 1.3 by employing the Witten deformation by a Morse function. Then we need to disassemble the trace of an operator on the Hilbert space of $L^{2}$-sections of a vector bundle over $M$ with respect to the cover (1.1). In Section 2, we introduce a piecewise trace-class operator and its piecewise trace which is the desired disassembly of the usual (or Roe’s) trace. In Section 3, we introduce an operator whose kernel function has propagation subordinate to a Gaussian function centered at the diagonal set, which we call an operator with Gaussian propagation. We prove the trace property of the piecewise trace of a product of operators with Gaussian propagation. We also consider the functional calculus of a generalized Dirac operators such that the resulting operator has Gaussian propagation. In Section 4, we introduce a disassembly of the normalized Betti number of Roe [19, 20], which is related to the $L^{2}$-Betti number of $M$, with respect to the cover (1.1), and prove its invariance under a certain deformation, including the Witten deformation by a bounded function. In Section 5, we collect all results obtained so far, and apply the Witten deformation to prove Theorem 1.3. Some conditions on a Morse function that we assume seem to be quite technical. In Section 6, we pose some questions on such technical conditions.
@@ 2.1. Fundamental domain
First of all, we construct a fundamental domain. Choose a $G$-invariant triangulation $L$ of $M$, that is, the lift of a triangulation of $M/G$. Then $L/G$ is a triangulation of $M/G$. For each open $n$-simplex of $L/G$, we choose one lift to $L$. For each open $k$-simplex with $k<n$, we choose one lift to $M$ which is in the closure of some open $n$-simplex of $M$ that we have chosen. Now we define a fundamental domain $K$ as the union of those open simplices of $M$. Observe that the closure of $K$ is the union of the closure of open $n$-simplices in $K$. Then the closure of $K$ is a finite complex, and a fundamental domain in the sense of [12]. Remark that a fundamental domain $K$ needs not be connected. By definition, we have
$M=\coprod_{g\in G}gK.$
In the sequel, we fix a fundamental domain $K$.
@@ 5.1. Strongly uniform Morse function
The key technical ingredient of Morse theory is the Morse lemma which states that near a critical point, a Morse function is identified with a nondegenerate quadratic function by choosing an appropriate coordinate. Such a coordinate is not generally a normal coordinate, which is inconvenient for our purpose. Then, instead of Morse lemma, we consider replacement of a given Morse function by a nice Morse function with the same critical points. We show that this is possible for a strongly uniform Morse function, and so we only need to prove Theorem 1.3 for such a nice Morse function. In the sequel, let $L$ denote the $G$-invariant triangulation of $M$ which is used to construct a fundamental domain $K$.
@@
A bounded Morse function $f\colon M\to\mathbb{R}$ is called strongly uniform if it is uniform for $\epsilon>0$ and satisfies the following conditions:
  1. (1)
    there is $\delta_{1}>0$ such that $|\lambda|\geq\delta_{1}$, where $\lambda$ ranges over all eigenvalues of the Hessian $\nabla^{2}f(p)$ for all $p\in\operatorname{Crit}(f)$;
  2. (2)
    there is $\delta_{2}>0$ such that $|\nabla f|\geq\delta_{2}$ on $M-N(\operatorname{Crit}(f);\epsilon_{0})$;
  3. (3)
    the $\epsilon$-neighborhood of any critical point of $f$ is contained in some $n$-simplex of $L$.
We set
@@
Let $g$ denote the fixed $G$-invariant metric on $M$. We introduce a new $G$-invariant metric on $M$. Take any small $\epsilon>0$. We define a $G$-invariant metric $g(\epsilon)$ on $M$ such that $g(\epsilon)=g$ on $N(L_{n-1};\frac{\epsilon}{2})$ and $g(\epsilon)$ is flat on the complement of $N(L_{n-1};\epsilon)$, where $L_{n-1}$ denotes the $(n-1)$-skeleton of $L$. Indeed, such a $G$-invariant metric $g(\epsilon)$ exists. Let $f\colon M\to\mathbb{R}$ be a uniform Morse function with respect to $\epsilon>0$. Then for any critical point $p$ of $f$, $N(p;\epsilon)$ is in the complement of $N(L_{n-1};\epsilon)$, and so the metric $g(\epsilon)$ is flat on $N(p;\epsilon)$. We remark that all results in the previous sections hold for any $G$-invariant metric on $M$, hence for $g(\epsilon)$. As mentioned above, boundedness of a function is independent of the choice of a $G$-invariant metric on $M$. We will use the fixed metric $g$ for considering upper and lower bounds for a Morse function.
@@
Take any $p\in\operatorname{Crit}_{k}(\tilde{f})$. Then $p\in g(\mathrm{Int}(K))$ for some $g\in G$, so $g$ is unique. Let $\rho_{p}(x)=\tilde{\rho}(d(x,p))$. Recall that we are working with the metric $g(\epsilon)$ on $M$, which is flat on $N(p;\epsilon)$ for each $p\in\operatorname{Crit}(\tilde{f})$. Let $\lambda_{1},\ldots,\lambda_{n}$ be the eigenvalues of the Hessian $\nabla^{2}\tilde{f}(p)$, and let
@@
As in the proof of Proposition 2.12, we have $\mu(\bar{\mathcal{I}})=0$. Then the statement follows from Theorem 1.3 and the positivity of $\mu$. ∎
@@
Then as in the proof of Corollary 1.4, we get $\mu(c_{k})\geq b_{k}^{(2)}$ for any $G$-invariant mean $\mu\colon\ell^{\infty}(G)\to\mathbb{R}$. Hence by Proposition 5.8, $c_{k}(g)\neq 0$ for infinitely many $g\in G$. ∎
@@
6. Questions
In this section, we pose questions concerning the conditions on a Morse function. boundedness and uniformness of a Morse function is irrelevant to the $G$-action on $M$, and the strong uniformness of a Morse functions also irrelevant to the $G$-action on $M$ except for the third condition in Definition 5.1. However, the relation of the third condition in Definition 5.1 and the $G$-action on $M$ is insignificant, and the authors believe that it is possible to drop this condition to prove Theorem 1.3. Then we pose:
Question 6.1.
Is it possible to prove Theorem 1.3 without the third condition in Definition 5.1?